At what launch angle or angles will a projectile land at 0.2 of its maximum possible range on flat ground? please show details
Answer #1

By R = u^2sin2?/g
=>R(max) = u^2/g [for? = 45*] -----------(i)
=>Let the angle is ? in 2nd condition:-
=>By R = u^2sin2?/g
=>0.2R(max) = u^2sin2?/g
=>0.2 x u^2/g = u^2sin2?/g
=>sin2? = 0.2

=>sin2? = sin12*

=>? = 6*